OpenStudy (anonymous):

Calculates the acceleration of a body starting from rest and has a speed of 20m / s after traveling 100m. Help please!

OpenStudy (anonymous):

a = v^2 \r answer=4 i think

OpenStudy (anonymous):

but above eq is for the body moving in circular path!

OpenStudy (anonymous):

ok assume constant acceleration using third equation \[2as=v^2 _{f}-vi^2_{i}\] here a=acceleration vf=final speed s=distance covered vi=initial velocity in the problem given data is s=100m vi=0 (bcz starting frm rest) vf=20m/s put in above equation \[a=\frac{v^2_{f}}{2s}\] \[a=\frac{20m^2}{2(100m)}\] i hope you can do this now

OpenStudy (anonymous):

\[a=\frac{(20m/s)^2}{2(100m)}\]

OpenStudy (anonymous):

@sami-21 this equation is for body moving uniformaly!

OpenStudy (anonymous):

x=xo+vo.t+1/2.a.t2 kia hum yeh wala formula use nhi kr sakhty @sami-21

OpenStudy (anonymous):

@ali110 look at first line of my response .

OpenStudy (anonymous):

muskan dear no because hamein time nahi given hai hamne sirf given info use karni hai

OpenStudy (anonymous):

@Muskan no because time is not given :)

OpenStudy (anonymous):

2as=v2f−vi2i yeh formula meri book may hai .. 2as?? means

OpenStudy (anonymous):

2 *aceeleration *distance

OpenStudy (anonymous):

nhi hai *****

OpenStudy (anonymous):

muskan y??

OpenStudy (anonymous):

2as is simply product of acceleration and distance it is in the standard formula . does not represent any physical Quantity.

OpenStudy (anonymous):

{s} = displacement {u} = initial velocity {v} = final velocity {a} = uniform acceleration

OpenStudy (anonymous):

okkk... Thanks Friends